OpenStudy (anonymous):

The numbers 20,21,29 form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? A) 21,22,30 B)40,42,58 C)60,66,87 D)70,84,116 Answer/Explanation PLEASE!! :-)

OpenStudy (shubhamsrg):

what do you mean by a pythagorean triplet ?

Parth (parthkohli):

The triplets are written in the form \(a,b,c\). The only work you have to do is to check if \(a^2 + b^2 = c^2\). Okay?

OpenStudy (shubhamsrg):

well actually you dont have to check by that method,,it'll take more time..there's another shortcut which i shall tell if she follows me up..

OpenStudy (shubhamsrg):

i wont say shortcut,,its actually quite obvious method,,

Parth (parthkohli):

Yep. I know that little shortcut ;)

Parth (parthkohli):

Well. I just don't want to confuse the asker, so I am going pretty basic.

OpenStudy (shubhamsrg):

hmm.. :|

OpenStudy (anonymous):

I prefer Parth's explanation. A Pythagorean triple follows the rule that squaring the two smaller numbers and adding them up results in the square of the larger number. Or in other words, a^2 + b^2 = c^2

OpenStudy (shubhamsrg):

so @dumbo123 ,,what can you infer from 20,21 and 29 being pythagoras triplets?

OpenStudy (shubhamsrg):

yep..method's hardly any problem.. i was just trying to make use of what info is given in the ques..

Parth (parthkohli):

Sometimes, there are those 'Can these be the sides of a right triangle?'... they are just the same.

OpenStudy (anonymous):

so it would be B! Thank you guys!

OpenStudy (anonymous):

Yes! Good job =)